Gold has a face centered cubic unit cell and has the density πŸπŸ—. πŸ‘ π’ˆ π’„π’Žβˆ’πŸ‘. Calculate the

length along an edge. (𝐴𝑒 = 197 𝑔 π‘šπ‘œπ‘™βˆ’1)


1
Expert's answer
2021-11-21T17:16:15-0500

FCC type gold contain atom

"\frac{1}{8}\times8+6\times\frac{1}{2}=4"

Mass of unit cell of FCC type

"4\times\frac{197}{6.023\times10^{23}}=1.30\times10^{-21}" g

Density of gold =19.3g/cm3

Density="\frac{mass of unit cell }{volume of unit cell}"

Volume of unit cell ="\frac{mass of unit cell }{density of gold}"


"V=\frac{1.30\times10^{-21}}{19.3}=6.73\times10^{-23}cm^3"

"V=a^3=6.73\times10^{-23}cm^3"

"a=4.06\times10^{-8}cm"

For FCC type unit cell

"a=\sqrt{8}r"


"r=\frac{a}{\sqrt8}=\frac{4.06\times10^{-8}}{\sqrt8}=1.43\times10^{-8}cm"